OpenStudy (anonymous):

The percentage of automobile cunsumers who are under 50 years of age degreased linerly from 55.9% in 1990 to 52.4% in 2005. Predeict when the percentage will be 48%?

hero (hero):

Just use the slope formula to get the slope, then use the point slope formula to find the equation

OpenStudy (anonymous):

So, \[52.4-55.9/105-90\]?

hero (hero):

yes, that's a good start.

OpenStudy (anonymous):

Well somewhere I'm getting hung up...so i end up with -3.5/15. Then go \[55.9=-3.5/15(90)+b? \]

OpenStudy (anonymous):

if this is right, that's where I'm stuck. I now have to predict when it will be 48%? Plug 48 into y? Solve

hero (hero):

Wait a sec...

OpenStudy (anonymous):

K :S

hero (hero):

Sorry, I thought you were using the point slope form. You decided to do this the other way... What do you get for b?

OpenStudy (anonymous):

Well originally 76.9

hero (hero):

That's correct

hero (hero):

So what is your y = mx + b formula now?

OpenStudy (anonymous):

Should i plug the 48 in as y?

hero (hero):

yes

OpenStudy (anonymous):

\[48=-3.5/15x+76.9\] Solve for x

hero (hero):

yes,and what do you get for x?

OpenStudy (anonymous):

-26.57?

hero (hero):

I get year 2023 and 10.2/12 months

OpenStudy (anonymous):

Hmm...so first I -76.9 from both sides?

hero (hero):

so right around October of year 2023

hero (hero):

Yes correct

OpenStudy (anonymous):

End up with -28.9 = -3.5/15x

hero (hero):

yes

OpenStudy (anonymous):

Do I convert m to -2.33?

hero (hero):

wait a sec...

hero (hero):

You should just solve it without converting.

OpenStudy (anonymous):

So did you get 433.5/3.5?

OpenStudy (anonymous):

123.86

OpenStudy (anonymous):

So in years 1900 + 123.6 = 2023.6? That seems a lot close to what you had

hero (hero):

Close enough. Good job

OpenStudy (anonymous):

Wow...what a pain. OK, well thank you for the time. That is a lot closer to a realistic answer. Just need to make sure I follow each step carefully. Thanks again.